OpenStudy (anonymous):

What values of x make the two expression below equal? (2x+1)(x-7)/11(x-7)=2x+1/11 A.) All real numbers expect -1/2 B.) All real numbers C.) All real numbers expect -1/2 and 7 D.) All real numbers expect 7 Please HELP! @saifoo.khan @iambatman @mathgeek27 @Joel_the_boss @mathmath333 @ganeshie8 @undeadknight26

OpenStudy (saifoo.khan):

What do you think?

OpenStudy (anonymous):

Um I really don't know do you plug in something?

OpenStudy (saifoo.khan):

No. Do you know about restrictions?

OpenStudy (anonymous):

No I do not.

OpenStudy (saifoo.khan):

\[\frac{(2x+1)(x-7)}{11(x-7)}=\frac{2x+1}{11}\] Okay, so if I plug x =7, it will be: \[\frac{(2x+1)(x-7)}{11(7-7)}=\frac{2x+1}{11}\] \[\frac{(2x+1)(x-7)}{0}=\frac{2x+1}{11}\] No fraction can be divided by 0 as it makes the fraction infinite. So x should NOT be equal to 7. This is called restriction.

OpenStudy (anonymous):

Okay so I would pick D?

OpenStudy (anonymous):

Okay, I don't understand what you just did.

OpenStudy (saifoo.khan):

Wait. it will be D. Ignore what I did last.

OpenStudy (anonymous):

Okay, thank you so much for your help!